# Parallel Collisionless Shocks in strongly Magnetized Electron-Ion Plasma. I. Temperature anisotropies

### v1

@@ Abstract
Collisionless electron-ion shocks are fundamental to astrophysical plasmas, yet their behavior in strong magnetic fields remains poorly understood. Using Particle-in-Cell (PIC) simulations with the SHARP-1D3V code, we investigate the role of the ion magnetization parameter $\sigma_{i}$ in parallel shock transitions. Strongly magnetized converging flows ($\sigma_{i}>1$) exhibit lower density compression ratios ($R\sim 2$), smaller entropy jumps, and suppressed particle acceleration, while maintaining pressure anisotropy stability due to conserved perpendicular temperatures across the shock, alongside increased parallel temperatures. In contrast, weakly magnetized shocks drive downstream mirror and firehose instabilities due to ion temperature anisotropy, which are suppressed in strongly magnetized cases. Additionally, weakly magnetized shocks exhibit the onset of a supra-thermal population induced by shock-drift acceleration, with most of the upstream kinetic energy thermalized for both electrons and ions in the downstream region. Our results demonstrate that perpendicular temperatures for both species are conserved in strongly magnetized cases and highlight deviations from standard ideal magnetohydrodynamic (MHD) behavior. These findings provide critical insights into the role of magnetic fields in parallel collisionless astrophysical shocks.
@@ 1 Introduction
Shock waves in fluids are fundamental phenomena that have been studied for over 100 years (Salas, 2007). In astrophysical context, shock waves are often collisionless, i.e. they propagate in a medium where the mean free path is much greater than the shock front. Since in a fluid, binary collisions mediate the transition between the upstream and the downstream, the absence of such collisions long cast some doubts on the very existence of collisionless shock waves (Sagdeev & Kennel, 1991), the study of which dates back to the 1960s with the pioneering work of Sagdeev (Sagdeev, 1966). Their reality has now been established for decades thanks, for example, to in situ observations of the bow shock of the Earth’s magnetosphere in the solar wind (Bale et al., 2003; Schwartz et al., 2011).
[…]
@@
In general, two potential sources of divergence between the behavior of a collisionless shock and that of an MHD shock, have been identified in recent years (see Bret (2020) and references therein). First, in the absence of collisions, a magnetized plasma can be stable even if anisotropic (Gary, 1993). Yet, MHD jump equations consider, at least in their simplest version, an isotropic medium. Secondly, it is known that collisionless shocks can accelerate particles (Axford et al., 1977; Blandford & Eichler, 1987) This is even the reason for the keen interest they arouse in astrophysics, since they are likely to explain, at least in part, the origin of cosmic rays. However, these accelerated particles escape the so to speak “Rankine Hugoniot budget” on which the MHD jump conditions are based (Berezhko & Ellison, 1999; David et al., 2022). This budget assumes that all of the upstream fluid passes into the downstream, together with the matter, momentum, and energy it carries. If some of it, the cosmic rays, moves back and forth around the shock front or returns from the downstream to the upstream, this assumption of integral transport from the downstream to the upstream is no longer valid.
@@
In the case of a parallel collisionless shock, i.e. a shock that propagates parallel to the ambient magnetic fields, Bret & Narayan (2018) developed a model capable of accounting for these phenomena. This model considered the simple case of a shock in a pair plasma, which makes it possible to treat the parallel and perpendicular temperatures of electrons and positrons on an equal footing. Confirmed by Particle-in-Cell (PIC) simulations (Haggerty et al., 2022), the main conclusion is that a strong parallel magnetic field reduce the density ratio $R$ of a strong sonic shock from 4 to 2. Such a reduction in the density jump would have significant effects on the production of cosmic rays since the power-law index of their momentum spectrum is $3R/(R-1)$ (Blandford & Ostriker, 1978).
Further studies varying field obliquity showed that the departure from MHD is indeed maximum for parallel shocks and minimum for perpendicular shocks (Bret & Narayan, 2019; Bret et al., 2024). It is therefore relevant to emphasize parallel shocks. In particular, while the model of Bret & Narayan (2018) and the PIC simulations confirming it (Haggerty et al., 2022), considered pair plasmas, it is important to assess the case of electrons/ions plasmas. Such is the aim of the present article.
@@
Pristine solar wind has been observed to exhibit increasing temperature anisotropy closer to the Sun. Initial measurements of proton temperature anisotropy by Wind at $1$ au (Bale et al., 2009) were later confirmed by Parker Solar Probe (PSP) (Huang et al., 2020), which revealed even more extreme anisotropy at smaller heliocentric distances (i.e., lower plasma beta), at $<0.25$ au. To distinguish shock-induced temperature anisotropy from instabilities inherent to the quiescent solar wind, we perform a kinetic-scale analysis of the former for a parallel shock propagating in a laminar medium. However, the transition layer rapidly develops magnetic field fluctuations, leading to local deviations from the parallel geometry, as observed by PSP (Jebaraj et al., 2024). A significant consequence of these fluctuations is their impact on particle acceleration, as demonstrated by test particle simulations (Fraschetti & Giacalone, 2015).
@@ 3 Particle-in-cell simulations
Figure 1: The spatial distribution of the number density (electron+ions) in various simulations at $t\omega_{i}=9400$. This shows that for values of $\sigma_{i}^{\prime}>4$ the compression ratio $R=n/n_{0}$ drops to $R\sim 2$ mean while it is $R>2$ for smaller values of $\sigma_{i}^{\prime}$. Dashed lines for each simulation represent the function in Equation  (8) using the best fit values of $R,\lambda_{\rm sh}$, and $x_{\rm sh}$ to number density profile. When plotting the density profile in simulations (solid lines) we filter all fluctuations on scales smaller than $4c/\omega_{i}$ using a moving average filter. The gray-shaded regions, which extend over about $115\,c/\omega_{i}$, define the regions where we compute the numerical entropy density in simulations.
@@
We employ a realistic ion-to-electron mass ratio, $m_{r}\equiv m_{i}/m_{e}=1836$, where $m_{i}$ ($m_{e}$) is the ion (electron) rest mass. Additionally, we include a parallel large-scale magnetic field $B_{0}$, and all simulations are conducted in the downstream/contact discontinuity rest frame, following the setup used in (Shalaby et al., 2022). While it is common practice to use a reduced ion-to-electron mass ratio in PIC simulations, Shalaby (2024) has shown that a reduced mass ratio can significantly influence both thermal and non-thermal energy dissipation in parallel shocks across a range of Alfvénic Mach numbers. A reduced mass ratio can significantly alter the hierarchy of the instabilities involved in the shock formation (Bret & Dieckmann, 2010). Therefore, we adopt the realistic mass ratio $m_{r}=1836$ in this work.
@@
The initial conditions for particles are identical in all simulations. Electrons and ions are initially in thermal equilibrium with $k_{B}T_{e}=k_{B}T_{i}=3.67\times 10^{-5}m_{e}c^{2}$ with isotropic temperatures, i.e., $k_{B}T_{x}=k_{B}T_{y}=k_{B}T_{z}=1.224\times 10^{-5}m_{e}c^{2}$. This fixes the upstream sonic Mach number to $\mathcal{M}_{\rm s}=387.3\times R/(R-1)$. Both electrons and ions in the far-upstream region drift toward the contact discontinuity at $x=0$ with speed $v_{u}=-0.1c$. We note that the temperatures $T_{i}$ and $T_{e}$ match measured plasma values in the solar wind at $1$ au (see, e.g., Maruca et al., 2023), whereas other plasma parameters, e.g., $n_{0}$ and $B_{0}$, or the shock speed exceed them by a few orders of magnitude, due to computational limitation. However, the $\mathcal{M}_{A}$ range covered here (and therefore the fast magnetosonic number) matches the supercritical regime of IPs at $1$ au, making the simulations relevant to the interpretation of observations.
@@
$\beta=\dfrac{1.333\times 10^{-6}}{\sigma_{i}^{{}^{\prime}}}$ (7)
@@
In the following sections, we only display the domain up to $1600\,c/\omega_{i}$ to focus on the shock transition regions.
@@ 4 Simulation Results
In this section, we present the results of our simulations. We examine how the density jump, entropy jump, temperature jump, and particle acceleration (for both electrons and ions) vary as the system transitions to the strongly magnetized regime i.e., $\sigma_{i}\geq 1$.
@@ 4.1 Number density across shock transition
The total number density (for electrons and ions combined) profile for various simulations at $t\omega_{i}=9400$ is shown in Figure 1. To characterize the density jump at various shocks, we fit the total number density across the shock with the following fit function (following Giacalone & Jokipii (2009) and as done by Haggerty et al. (2022))
@@
This demonstrates that the compression ratio decreases to the expected value of $R\sim 2$ for strongly magnetized simulations ($\sigma_{i}>1$). At low magnetization, particularly for $\sigma_{i}^{\prime}<1$, the downstream density exhibits large amplitude fluctuations (see the shaded region in Figure 1), indicating that the downstream plasma has not reached a steady state at $t\omega_{i}=9400$. This results in best-fit compression ratios for $R$ significantly larger than $4$. At later times, when an approximate steady state is achieved, a higher compressibility ($R>4$) is reached in simulations due to the presence of accelerated particles (Blondin & Ellison, 2001; Fraschetti et al., 2010), which is also seen in kinetic scale simulations  (Shalaby et al., 2022; Gupta et al., 2024).
@@ 4.2 Entropy across shock transition
In this section, we study the change in entropy across various simulated shocks. To compute the entropy jump, we need to approximate the numerical values for the entropy.
If we do not assume that the distribution in momenta is separable in each direction but rather depends only on the energy (similar to the case of an isotropic Maxwell-Jüttner distribution in the rest frame), we can proceed as follows. If we divide the particle energy/momenta into energy/momenta bins labeled with $k$ and of equal size $\Delta u$, we can write:
@@
where $N_{k}$ is the number of particles with energies within the $k$th energy bin, and $N_{p}$ is the total number of particles.
@@
$\Rightarrow\frac{s}{k_{B}}\rightarrow\sum_{k}4\pi u_{k}^{2}\Delta u\,f_{k}\ln f_{k}=\sum_{k}N_{k}\ln\frac{N_{k}}{4\pi u_{k}^{2}\Delta u}.\quad$ (10)
@@
where, $u^{j}_{\perp}$ is the value of bin centers in the perpendicular direction. To compute the entropy jump across various shocks, we calculate the entropy in the downstream and upstream regions, i.e., the gray-shaded regions in Figure 1, which extend over about $115\,c/\omega_{i}$ and use Equation (13) to compute the entropy densities numerically. The entropy jump is then defined as $\Delta s=s_{\rm downstream}-s_{\rm upstream}$ for various simulations. The dependence of $\Delta s$ on the value of $\sigma_{i}$ is shown in the bottom panel of Figure 2 for both ions (red dashed) and electrons (black dashed) in various simulations.
For both electron and ion species, as shown in Figure 2, an entropy jump is found in all simulations where the value of such jump is typically smaller in the strongly magnetized simulations with $\sigma_{i}\geq 1$.
@@ 4.3 Phase-space particle distribution across shocks
In this section, we present the phase-space distribution of particles (electrons in Figure 3 and ions in Figure 4) across various shocks at $t\omega_{i}=9400$. In both figures, we plot the 2D phase-space distribution for parallel momentum, i.e., $f(u_{x},x)$, in the left column, where $u_{j}=v_{j}\gamma$ is the $j$ component of the particle 4-velocity, while the perpendicular momentum distribution, $f(u_{y},x)$, are shown in the right column.
@@
In all simulations, the upstream kinetic energy of electrons is fully thermalized in the downstream region. Parallel and quasi-parallel shocks that satisfy $21<\mathcal{M}_{A}\lesssim\sqrt{m_{r}/R}/$ 2 can efficiently accelerate electrons (Shalaby et al., 2022). This is due to the destabilization of intermediate-scale unstable modes in these shocks (Shalaby et al., 2021, 2023). Indeed, for such shocks (top two rows in Figure 3), electrons are clearly accelerated and escape toward the upstream region of the shock. For the simulations with $\sigma_{i}^{\prime}=4$ and $10$, electron acceleration is absent and full electron thermalization in the downstream region has already occurred at the time $t\omega_{i}=9400$.
@@
A large velocity component along the shock surface, $u_{y}$, comparable to $u_{x}$, is observed in the two simulations with $\sigma<1$ for both ions and electrons (top two rows of Figures 3 and 4). These two populations are likely to be efficiently energized into the supra-thermal range by shock drift acceleration via the motional electric field $|v_{\rm sh}\times B|$ (see the bottom panel in Figure 6): a large upstream $B_{\perp}$ component becomes dominant over $B_{\parallel}$ in the downstream plasma. Since the oblique magnetic field persists past the shock ramp for hundreds of $c/\omega_{pi}$, the population with large $u_{y}$ homogeneously fills the downstream region over a comparable length.
@@
An important consequence of this finding is that the supra-thermal populations of ions and electrons are driven by magnetic fluctuations induced by temperature anisotropy at the shock, rather than by fluctuations pre-existing to the shock passage or by the streaming of resonant energized ions (which have not yet been generated). An in-depth analysis of this effect will be presented in a companion paper. For shocks with $\sigma_{i}<1$, while some ions and electrons become supra-thermal, most of the upstream kinetic energy is thermalized in the downstream region.
@@
For simulations with $\sigma_{i}^{\prime}=4$ and $10$, at time $t\omega_{i}=9400$, thermalization of ions in the downstream region has not yet fully occurred. Nevertheless, the visible two-stream distribution in the downstream region are hotter compared to that of the incoming ion stream. As we will see in the next section, the parallel temperature, in fact, increases by a factor of 100 (black dashed curve in the top panel of Figure 5) in the downstream region compared to that of the far upstream. In this work, we adopt the convention of referring to the denser part of the domain as the downstream region, regardless of whether it qualifies as a shock. This classification does not require the parallel ion momentum to be fully thermalized in the downstream region. Notably, in all simulations, the electrons are already fully thermalized in the downstream region.

### v2

@@ Abstract
Collisionless electron-ion shocks are fundamental to astrophysical plasmas, yet their behavior in strong magnetic fields remains poorly understood. Using Particle-in-Cell (PIC) simulations with the SHARP-1D3V code, we investigate the role of the ion magnetization parameter $\sigma_{i}$ in parallel shock transitions. Strongly magnetized converging flows ($\sigma_{i}>1$) exhibit lower density compression ratios ($R\sim 2$), smaller entropy jumps, and suppressed particle acceleration, while maintaining pressure anisotropy stability due to conserved perpendicular temperatures across the transition region, alongside increased parallel temperatures. In contrast, weakly magnetized shocks drive downstream mirror and firehose instabilities due to ion temperature anisotropy, which are suppressed in strongly magnetized cases. Additionally, weakly magnetized shocks exhibit the onset of a supra-thermal population induced by shock-drift acceleration, with most of the upstream kinetic energy thermalized for both electrons and ions in the downstream region. Our results demonstrate that perpendicular temperatures for both species are conserved in weakly and strongly magnetized cases and highlight deviations from standard ideal magnetohydrodynamic (MHD) behavior in strongly magnetized cases. These findings provide critical insights into the role of magnetic fields in parallel collisionless astrophysical shocks.
@@ 1 Introduction
Shock waves in fluids are fundamental phenomena that have been studied for over 100 years (Tidman & Krall, 1971; Whitham, 1974; Zel’dovich & Raizer, 2002; Landau & Lifshitz, 2013; Salas, 2007). In astrophysical context, shock waves are often collisionless, i.e. they propagate in a medium where the mean free path is much greater than the shock front thickness (Balogh & Treumann, 2013; Burgess & Scholer, 2015). Since in a fluid, binary collisions mediate the transition between the upstream and the downstream, the absence of such collisions long cast some doubts on the very existence of collisionless shock waves (Sagdeev & Kennel, 1991), the study of which dates back to the 1960s with the pioneering work of Sagdeev (Sagdeev, 1966). Their reality has now been established for decades thanks, for example, to in situ observations of the bow shock of the Earth’s magnetosphere in the solar wind (Bale et al., 2003; Schwartz et al., 2011).
It is now well-established that shocks are ubiquitous in astrophysics, ranging from interplanetary shocks in the solar system to supernova remnants (McKee & Hollenbach, 1980). In-situ measurements at Earth’s bow shock have revealed temperature anisotropies in the downstream plasma, which are known to trigger firehose instabilities at quasi-parallel shocks and mirror instabilities at quasi-perpendicular shocks (Bale et al., 2009; Maruca et al., 2011; Schlickeiser et al., 2011). For instance, Wind measurements (Vogl et al., 2001) during an oblique crossing of Earth’s bow shock show $\sim\%1$ deviations in density compression, which were explained therein by pressure anisotropies. Similarly, Voyager 1 data from the quasi-perpendicular solar wind termination shock crossing provided evidence of mirror instabilities in downstream region (Liu et al., 2007; Génot, 2008).
@@
In general, two potential sources of divergence between the behavior of a collisionless shock and that of an MHD shock, have been identified in recent years (see Bret (2020) and references therein). First, in the absence of collisions, a magnetized plasma can be stable even if anisotropic (Gary, 1993). Yet, MHD jump equations consider, at least in their simplest version, an isotropic medium. Secondly, it is known that collisionless shocks can accelerate particles (Axford et al., 1977; Blandford & Eichler, 1987), which is the primary reason for their significant interest in astrophysics, as they may partly explain the origin of cosmic rays.
However, these accelerated particles escape the so to speak “Rankine Hugoniot budget” on which the MHD jump conditions are based (Berezhko & Ellison, 1999; David et al., 2022). This budget assumes that all of the upstream fluid passes into the downstream, together with the matter, momentum, and energy it carries. If some of it, the cosmic rays, moves back and forth around the shock front or returns from the downstream to the upstream, this assumption of integral transport from the downstream to the upstream is no longer valid.
@@
In the case of a parallel collisionless shock, i.e. a shock that propagates parallel to the ambient magnetic fields, Bret & Narayan (2018) developed a model capable of accounting for these phenomena. This model considered the simple case of a shock in a pair plasma, which makes it possible to treat the parallel and perpendicular temperatures of electrons and positrons on an equal footing. Confirmed by Particle-in-Cell (PIC) simulations (Haggerty et al., 2022), the main conclusion is that a strong parallel magnetic field reduce the density ratio $R$ of a strong sonic shock from 4 to 2. Such a reduction in the density jump would have significant effects on the production of cosmic rays since the power-law index of their momentum spectrum is 3R/(R1)3R/(R-1) (Blandford & Ostriker, 1978).
Further studies varying field obliquity showed that the departure from MHD is indeed maximum for parallel shocks and minimum for perpendicular shocks (Bret & Narayan, 2019; Bret et al., 2024). It is therefore relevant to emphasize parallel shocks. In particular, while the model of Bret & Narayan (2018) and the PIC simulations confirming it (Haggerty et al., 2022) considered pair plasmas, it is important to assess the case of electrons/ions plasmas. Such is the aim of the present article.
@@
Pristine solar wind has been observed to exhibit increasing temperature anisotropy closer to the Sun. Initial measurements of proton temperature anisotropy by Wind at $1$ au (Bale et al., 2009) were later confirmed by Parker Solar Probe (PSP) (Huang et al., 2020, 2025), which revealed even more extreme anisotropy at smaller heliocentric distances (i.e., lower plasma beta), at <0.25<0.25 au. To distinguish shock-induced temperature anisotropy from instabilities inherent to the quiescent solar wind, we perform a kinetic-scale analysis of the former for a parallel shock propagating in a laminar medium. However, the transition layer rapidly develops magnetic field fluctuations, leading to local deviations from the parallel geometry, as observed by PSP (e.g., Jebaraj et al., 2024). A significant consequence of these instability-related, but also pre-existing, fluctuations is their impact on particle acceleration, as demonstrated by test particle simulations (e.g., Fraschetti & Giacalone, 2015).
@@ 3 Particle-in-cell simulations
Figure 1: The spatial distribution of the number density (electron+ions) in various simulations at $t\omega_{i}=9400$. This shows that for values of $\sigma_{i}^{\prime}>4$ the compression ratio $R=n/n_{0}$ drops to $R\sim 2$ mean while it is $R>2$ for smaller values of $\sigma_{i}^{\prime}$. Dashed lines for each simulation represent the function in Equation  (8) using the best fit values of $R,\lambda_{\rm sh}$, and $x_{\rm sh}$ to number density profile. When plotting the density profile in simulations (solid lines) we filter all fluctuations on scales smaller than $4c/\omega_{i}$ using a moving average filter. The gray-shaded regions, which extend over about $115\,c/\omega_{i}$, define the regions where we compute the numerical entropy density in simulations (Figure 2, bottom panel).
@@
We employ a realistic ion-to-electron mass ratio, $m_{r}\equiv m_{i}/m_{e}=1836$, where $m_{i}$ ($m_{e}$) is the ion (electron) rest mass. Additionally, we include a parallel large-scale magnetic field $B_{0}$, and all simulations are conducted in the downstream/contact discontinuity rest frame, following the setup used in Shalaby et al. (2022). While it is common practice to use a reduced ion-to-electron mass ratio in PIC simulations, Shalaby (2024) has shown that a reduced mass ratio can significantly influence both thermal and non-thermal energy dissipation in parallel shocks across a range of Alfvénic Mach numbers. A reduced mass ratio can significantly alter the hierarchy of the instabilities involved in the shock formation (Bret & Dieckmann, 2010). Therefore, we adopt the realistic mass ratio $m_{r}=1836$ in this work.
@@
The initial conditions for particles are identical in all simulations. Electrons and ions are initially in thermal equilibrium with $k_{B}T_{e}=k_{B}T_{i}=3.67\times 10^{-5}m_{e}c^{2}$ with isotropic temperatures, i.e., $k_{B}T_{x}=k_{B}T_{y}=k_{B}T_{z}=1.224\times 10^{-5}m_{e}c^{2}$. This fixes the upstream sonic Mach number to $\mathcal{M}_{\rm s}=v_{\rm sh}/\sqrt{2\Gamma k_{B}T_{i}/m_{i}}=387.3\times R/(R-1)$, where $\Gamma=5/3$ is the adiabatic index. Both electrons and ions in the far-upstream region drift toward the contact discontinuity at $x=0$ with speed $v_{u}=-0.1c$. We note that the temperatures $T_{i}$ and $T_{e}$ match measured plasma values in the solar wind at $1$ au (see, e.g., Maruca et al., 2023), whereas other plasma parameters, e.g., $n_{0}$ and $B_{0}$, or the shock speed exceed them by a few orders of magnitude, due to computational limitation. However, the $\mathcal{M}_{A}$ range covered here (and therefore the fast magnetosonic number $M_{f}$) matches the supercritical regime of interplanetary shocks at $1$ au ($M_{f}\gtrsim 1.5$ in low plasma beta regime for parallel shocks), making the simulations relevant to the interpretation of observations.
@@
$\beta=\dfrac{1.333\times 10^{-6}}{\sigma_{i}^{{}^{\prime}}}.$ (7)
@@
In the following sections, we only display the domain up to $1600\,c/\omega_{i}$ from the reflecting wall to focus on the shock transition regions.
@@ 4 Simulation Results
In this section, we present the results of our simulations. We examine how the density jump, entropy jump, temperature jump, and particle acceleration (for both electrons and ions) vary as the system transitions to the strongly magnetized regime i.e., $\sigma_{i}\geq 1$. While we present the final stages of our simulations in this section, the full dynamical evolution for all simulations can be found at mohamadshalaby.github.io/Pshock_Sig01.html.
@@ 4.1 Number density across shock transition
The total plasma number density (for electrons and ions combined) profile for various simulations at $t\omega_{i}=9400$ is shown in Figure 1. To characterize the density jump at various shocks, we fit the total number density across the shock with the following fit function (following Giacalone & Jokipii (2009) and as done by Haggerty et al. (2022))
@@
This demonstrates that the compression ratio decreases to the expected value of $R\sim 2$ for strongly magnetized simulations ($\sigma_{i}>1$). At low magnetization, particularly for $\sigma_{i}^{\prime}<1$, the downstream density exhibits large amplitude fluctuations (see the shaded region in Figure 1), indicating that the downstream plasma has not reached a steady state at $t\omega_{i}=9400$. This results in best-fit compression ratios for $R$ significantly larger than $4$. At later times, when an approximate steady state is achieved, a higher compressibility ($R>4$) is reached in simulations due to the presence of accelerated particles (Blondin & Ellison, 2001; Fraschetti et al., 2010), which is also seen in other kinetic scale simulations  (Shalaby et al., 2022; Gupta et al., 2024).
@@ 4.2 Entropy across shock transition
In this section, we study the change in entropy across the transition region in various simulations. Below, we explain how we approximate the numerical values for the entropy.
If we do not assume that the distribution in momenta is separable in each direction but rather depends only on the energy (similar to the case of an isotropic Maxwell-Jüttner distribution in the rest frame), we can proceed as follows. If we divide the particle momenta, $u$, where $u=\gamma v$ is the spatial part of the 4-velocity, into momenta bins labeled with $k$ and of equal size $\Delta u$, we can write:
@@
where $N_{k}$ is the number of particles with energies within the $k$th energy bin, $N_{p}$ is the total number of particles, and all velocities are normalized by the speed of light, $c$.
@@
$\Rightarrow\frac{s}{k_{B}}\rightarrow\sum_{k}4\pi u_{k}^{2}\Delta u\,f_{k}\ln f_{k}=\sum_{k}N_{k}\ln\frac{N_{k}}{4\pi u_{k}^{2}\Delta u}.$
@@
where, $u^{j}_{\perp}$ is the value of bin centers in the perpendicular direction. To compute the entropy jump across the transition region in various simulations, we calculate the entropy in the downstream and upstream regions, i.e., the gray-shaded regions in Figure 1, which extend over about $115\,c/\omega_{i}$ and use Equation (13) to compute the entropy densities numerically. The entropy jump is then defined as $\Delta s=s_{\rm downstream}-s_{\rm upstream}$ for various simulations. The dependence of $\Delta s$ on the value of $\sigma_{i}$ is shown in the bottom panel of Figure 2 for both ions (red dashed) and electrons (black dashed) in various simulations.
For both electron and ion species, as shown in Figure 2, an entropy jump is found in all simulations where the value of such jump is typically smaller in the strongly magnetized simulations with $\sigma_{i}\geq 1$. However, as shown in the next section, for $\sigma_{i}^{\prime}=4$ and $10$, ions in the counter-propagating regions exhibit heating via electrostatic perturbations without achieving full thermalization. Consequently, the observed ion entropy jump in these cases stems solely from electrostatic heating rather than complete thermalization.
@@ 4.3 Phase-space particle distribution across shocks
In this section, we present the phase-space distribution of particles (electrons in Figure 3 and ions in Figure 4) in various simulations at $t\omega_{i}=9400$. In both figures, we plot the 2D phase-space distribution for parallel momentum, i.e., $f(u_{x},x)$, in the left column, where $u_{j}=v_{j}\gamma$ is the $j$ component of the particle 4-velocity, while the perpendicular momentum distribution, $f(u_{y},x)$, is shown in the right column.
In all simulations, the upstream kinetic energy of electrons is fully thermalized in the downstream region (see Figure 3). This thermalization results from the well-known two-stream instabilities of the initially counter-streaming electron-ion beams (Stix, 1992). These instabilities destabilize the parallel (i.e., along $B_{0}$) electrostatic wave modes both at electron and ion skin-depth scales [^1] When focusing on parallel modes, the linear dispersion relation splits into two parts: one related to electrostatic instabilities (independent of the large-scale magnetic field $B_{0}$), and another containing only electromagnetic wave modes. See Appendix B of (Shalaby et al., 2021) for details. In all simulations, these electric perturbations fully thermalize electrons in the beam interaction region on a very short time scale, $t\omega_{i}<10$. However, this is insufficient to completely thermalize ions (regardless of the $\sigma$ value) as we show below.
Notably, the separation of the electron-to-proton scale by using herein the realistic $m_{r}$ allowed to temporally separate the downstream thermalization of the two species (seen also for protons at $\sigma<1$). In addition, since the upstream ram pressure is proton-dominated, in such an early stage the dissipation into heating is insufficient to form a stationary shock transition.
@@
Parallel and quasi-parallel shocks that satisfy $1<\mathcal{M}_{A}\lesssim\sqrt{m_{r}/R}/2$ can efficiently accelerate electrons (Shalaby et al., 2022). This is due to the destabilization of intermediate-scale unstable modes in these shocks (Shalaby et al., 2021, 2023). Indeed, for such shocks (top two rows in Figure 3), electrons are clearly accelerated and escape toward the upstream region of the shock. For the simulations with $\sigma_{i}^{\prime}=4$ and $10$, electron acceleration is absent as expected.
@@
A large velocity component along the shock surface, $u_{y}$, comparable to $u_{x}$, is observed in the two simulations with $\sigma_{i}<1$ for both ions and electrons (top two rows of Figures 3 and 4). These two populations are likely to be efficiently energized into the supra-thermal range by shock drift acceleration via the motional electric field $|v_{\rm sh}\times B|$ (see section 4.5), which is expected to accelerate particles efficiently at quasi-perpendicular shocks (Jokipii, 1982; Decker, 1988; Krauss-Varban & Wu, 1989; Ball & Melrose, 2001; Fraschetti & Giacalone, 2015): a large upstream $B_{\perp}$ component becomes dominant over $B_{\parallel}$ in the downstream plasma. Since the oblique magnetic field persists past the shock ramp for hundreds of $c/\omega_{i}$, the electron population with large $u_{y}$ homogeneously fills the downstream region over a comparable length-scale.
@@
An important consequence of this finding is that the scattering of both ions and electrons into supra-thermal energies can be driven, or contributed, by magnetic fluctuations induced by temperature anisotropy at the shock (as pointed out by Quest & Shapiro, 1996), rather than by fluctuations pre-existing to the shock passage or by the streaming of resonant energized ions (which have not yet been generated). An in-depth analysis of this effect will be presented in a companion paper. For shocks with $\sigma_{i}<1$, while some ions and electrons become supra-thermal, most of the upstream kinetic energy is thermalized in the downstream region.
For simulations with $\sigma_{i}^{\prime}=4$ and $10$, at time $t\omega_{i}=9400$, the driven electrostatic instabilities are insufficient to thermalize ions in the interaction region, as shown in the bottom two panels of Figure 4. A similar configuration appears in the early stages of simulations with $\sigma_{i}^{\prime}<4$, where ion thermalization occurs afterward concurrently with the destabilization of parallel electromagnetic wave modes, as shown in the top two panels of Figure 4.
Solving the electromagnetic dispersion relation for counter-streaming electron-ion beams yields a stability cutoff of $\sigma_{i}^{\prime}=2$ for parallel electromagnetic wave modes (Quest, 1988). That is, parallel electromagnetic modes for counter-streaming ions become completely stable for $\sigma_{i}^{\prime}\geq 2$. This cutoff is independent of both the background (non-streaming) electron temperature and the counter-streaming ion plasma $\beta$ 222 To test this, we make use of the linear dispersion solver BO (Xie et al., 2021) which is publicly available at https://github.com/hsxie/bo. . For $\sigma_{i}^{\prime}<2$ simulations, the unstable electromagnetic wave modes not only thermalize ions in the downstream region but also facilitate the transfer of part of the upstream ion kinetic energy to electrons, as shown in the top two panels of Figure 3 and the top panel of Figure  5.
It is important to note that oblique electromagnetic wave modes may become unstable at resonant wavelengths, potentially enabling complete ion thermalization for $\sigma_{i}^{\prime}>2$. This possibility can be investigated through both linear analysis and ultimately 2D3V kinetic simulations; we defer this investigation to future work.
@@
For $\sigma_{i}^{\prime}>2$ cases, we will see in the next section that the parallel ion temperature, in fact, increases by a factor of 100 (black dashed curve in the top panel of Figure 5) in the downstream region compared to that of the far upstream due to these electrostatic unstable wavemodes. In this work, we adopt the convention of referring to the denser part of the domain as the downstream region, regardless of whether it qualifies as a shock. This classification does not require the parallel ion momentum to be fully thermalized in the downstream region. Notably, in all simulations, the electrons are already fully thermalized in the downstream region.